OpenStudy (anonymous):

What is the solution to the equation sqrt 2x+8-6=4 . Show your work.

OpenStudy (anonymous):

@Directrix

OpenStudy (bibby):

\(\sqrt {2x+8}-6=4\) ??

OpenStudy (anonymous):

my answer was x=46 but my teacher said my work is in complete soo and yes ths right

OpenStudy (anonymous):

@bibby

OpenStudy (bibby):

start by isolating the term holding the x in t his case it's \(\color{red}{\sqrt {2x+8}}-6=4\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so it would be sqrt 2x+8 right

OpenStudy (bibby):

what is "it"?

OpenStudy (anonymous):

X=2 i think im not great at math but i will try

OpenStudy (bibby):

I have no idea what you're doing start by getting the red term alone

OpenStudy (anonymous):

i have no idea on how to do that

OpenStudy (anonymous):

@KamiBug @iGreen

OpenStudy (kamibug):

46 is correct. Here's the work: Add 6 to both sides. sq.rt. 2x + 8 = 10 Square both sides. 2x + 8 = 100 Subtract 8 from both sides. 2x = 92 Divide both sides by 2. x = 46

OpenStudy (bibby):

spoonfeeding answers, lol

OpenStudy (anonymous):

thnak yall so much